# multi-sr.bib

### v1

@@ 3 Eigenvalues of smooth kernel matrices in the flat limit
The bounds of Theorem 1 are not necessarily tight. As we show in the next lemma, the exactness of the asymptotic orders depends on the analytic properties of the kernel $\mathcal{K}$ and the geometry of ${\cal X}$, therefore additional assumptions ought to be made on both of those objects in order to obtain matching lower bounds.
Lemma 1.
For $\mathcal{K}\in\mathcal{C^{(\infty,\infty)}}(\Omega)$ and ${\cal X}\subset\mathbb{R}^{d}$, there exists $C=C({\cal X},\mathcal{K})\in\mathbb{R}$ such that
In particular, $C=\det\left(\tilde{R}W\tilde{R}^{T}\right)$ where $\tilde{R}\in\mathbb{R}^{n\times p_{m}}$ is full rank, and $W=W^{\mathcal{K}}_{\leq m}\in\mathbb{R}^{p_{m}\times p_{m}}$ is the Wronskian matrix of $\mathcal{K}$ as in Definition 4.
Proof.
Consider the matrices $Q_{\perp,k-1}$ defined in the proof of Theorem 1. Let $\mathcal{M}_{k}\subseteq range(Q_{\perp,k-1})$ and $\mathcal{M}_{k}\perp range(Q_{\perp,k})$ where $\mathcal{M}_{k}\in\mathbb{R}^{n\times t_{k}}$ and $Q_{\perp,-1}=\mathbb{R}^{n\times n}$. Define
$\mathcal{Q}_{\perp}:=\big{(}\mathcal{M}_{0},\mathcal{M}_{1},...,\mathcal{M}_{m}\big{)}\in\mathbb{R}^{n\times n}$
It is easy to see that $\mathcal{Q}_{\perp}^{T}V_{\leq m}$ is upper triangluar, with the block-diagonal part given by
@@
Therefore by [usevich2021, Lemma 6.4] we have
@@
where the last equality follows from $\epsilon^{2{(m+1)}}E_{n}^{-1}=O(\epsilon)$ and $\tilde{R}W\tilde{R}^{T}\in\mathbb{R}^{n\times n}$ This implies that $\epsilon^{-2\sum_{j=0}^{m}jt_{j}}\det(K_{\epsilon})=\det(\tilde{R}W\tilde{R}^{T}+O(\epsilon))=\det(\tilde{R}W\tilde{R}^{T})+O(\epsilon)$ (by Lemma (8) in the Appendix) with $C=\det(\tilde{R}W\tilde{R}^{T})$.
Corollary 1.
Suppose $\mathcal{K}$ is symmetric and analytic (in both variables) in the neighborhood of $(\mathbf{0},\mathbf{0})$. The scaling of the eigenvalues as given in Theorem 1 is exact, if and only if $\det(\tilde{R}W\tilde{R}^{T})\neq 0$.
@@
By Theorem 1 and [usevich2021, Theorem 2.9], the eigenvalues of $K_{\epsilon}({\cal X})$ satisfy
@@
where $t_{k}:=\operatorname{rank}(V_{\leq k})-\operatorname{rank}(V_{\leq k-1})$ and $m=\mu({\cal X})$, and $\tilde{\lambda}_{s,j}$ do not depend on $\epsilon$ (but may be zero). By Lemma 1 we have that
@@
Now we just compare the two expressions above. In one direction, if the scaling is exact then $\tilde{\lambda}_{s,j}\neq 0$ for all $0\leq s\leq m$ and $0\leq j\leq t_{s}$, which implies $\left|\tilde{R}W\tilde{R}^{T}\right|\neq 0$. In the other direction, if there are $\ell\geq 1$ zeros $\tilde{\lambda}_{k}\in\{\tilde{\lambda}_{s,j}\}_{s=1,\dots,m}^{j=1,\dots,t_{m}}$ and $\tilde{\lambda}_{k}=0$ for $k=1,...,\ell$, then $\big{|}K_{\epsilon}\big{|}=\epsilon^{\ell+2\sum_{i=0}^{m}it_{i}}(\prod_{k=\ell+1}^{n}\tilde{\lambda}_{k}+O(\epsilon))$, forcing $\left|\tilde{R}W\tilde{R}^{T}\right|=0$. ∎
@@ 4 Eigenvalues of the multidimensional Dirichlet kernel
In this section we consider the specific case of the multidimensional Dirichlet kernel, which is directly motivated by the problem of super-resolution as outlined in the Introduction. Let ${\cal S}\subset\mathbb{R}^{d}$ denote a finite set of frequencies, symmetric around the origin, where the spectral data (2) is available. We first derive a generic sufficient condition on ${\cal S}$ which ensures that the scaling order of Theorem 1 is exact. Later we show that this condition is satisfied by ${\cal S}={\cal G}_{N}$, the regular symmetric grid in $\mathbb{R}^{d}$ of side length $12N+$ 1, for large enough $N$.
Definition 6.
@@ 4.1 A nondegeneracy condition
Note that $\mathcal{D}_{{\cal S}}$ is a translation invariant, symmetric and analytic kernel. Recall the definition (8) of the Wronskian matrix for an arbitrary kernel. Set $W:=W^{\mathcal{D_{S}}}_{\leq m}$. By Corollary 1, in order to show tightness, it is sufficient to demonstrate that $C:=\det(\tilde{R}W\tilde{R}^{T})\neq 0$. In the following lemma, we show that this holds under a certain (more or less natural) non-degeneracy assumption on the set ${\cal S}$.
Lemma 2.
For $W:=W_{\leq m}$ as defined above, $\det(\tilde{R}W\tilde{R}^{T})>0$ holds whenever $\operatorname{rank}(V_{\leq m}({\cal S}))\geq n$.
Remark 2.
@@
where $V_{\leq m}({\cal S})=\big{[}{(\omega)}^{\alpha}\big{]}_{\omega\in{\cal S}}^{\alpha\in\mathbb{P}_{m}}$ is the multivariate Vandermonde matrix and $F:=\operatorname{diag}\left(\frac{i^{|\alpha|}}{\alpha!}\right)_{\alpha\in\mathbb{P}_{m}}$. Let $B:=FV^{T}_{\leq m}({\cal S})$, and consider $B=U\Sigma T^{*}$ the singular value decomposition of $B$. We therefore have $W=BB^{*}=U\Sigma^{2}U^{*}$, implying that $W$ is positive semidefinite (since $W$ is Hermitian and has non-negative eigenvalues). Since $\tilde{R}$ is full rank by Lemma 1, it follows that $\tilde{R}W\tilde{R}^{T}$ is positive semidefinite as well, implying that $\det(\tilde{R}W\tilde{R}^{T})\geq 0$. To ensure $\det(\tilde{R}W\tilde{R}^{T})>0$, we require that $\tilde{R}W\tilde{R}^{T}$ be positive definite. The latter condition is satisfied when $\operatorname{rank}(\tilde{R}W\tilde{R}^{T})\geq n$, which is equivalent to $\operatorname{rank}(W)=\operatorname{rank}(V_{\leq m}({\cal S}))\geq n$. ∎
Remark 3.
For $d=1$, we have that $m=n-1$ and $p_{n-1}={n\choose 1}=n$, meaning that the condition $\operatorname{rank}(V_{\leq n-1}({\cal S}))\geq n$ in Lemma (2) is satisfied as long as the sampling set ${\cal S}$ contains at least $n$ distinct points.
Remark 4.
Given the condition $\operatorname{rank}(V_{\leq m}({\cal S}))\geq n$ in Lemma (2), we aim to find a sampling set ${\cal S}_{*}$ with $p_{\ell}$ points where $p_{\ell-1}<n$ and $p_{\ell}\geq n$ such that $\operatorname{rank}(V_{\leq\ell}({\cal S}_{*}))\geq n$ which is independent of $m=\mu({\cal X})$ (the geometry of the nodes).
By applying Corollary 1, we immediately obtain the following result.
Lemma 3.
@@ 4.2 Geometric characterization condition and the uniform grid
The rank condition in Lemma 2 is nontrivial to verify for a given set ${\cal S}$. In this section we provide an explicit example of such a set, namely, the uniform grid ${\cal G}_{N}$ in $\mathbb{R}^{d}$. Our approach is a straightforward application of well-known results in multivariate polynomial interpolation, adapted to our setting.
We first recall some definitions and results from [sauer2006].
Definition 7 ([sauer2006], p.194).
Let $\mathcal{P}$ be a linear subspace of the polynomial ring $\Pi:=\mathbb{R}\left[x_{1},...,x_{d}\right]$. The polynomial interpolation problem with respect to a (finite) set of linearly independent functionals $\Theta\subset\Pi^{\prime}$ ($\Pi^{\prime}$ is the dual space) is said to be poised for $\mathcal{P}$, if for any $Y\in\mathbb{R}^{\#\Theta}$ there exists a unique $f\in\mathcal{P}$ such that $\Theta f=Y$.
Theorem 2 ([sauer2006] p.194).
@@
For $\Theta_{{\cal S}}:=\{\delta_{\xi}\}_{\xi\in{\cal S}}$ the set of point evaluation functionals on ${\cal S}$, and the standard monomial basis $P_{n}$ of polynomials of total degree $\leq n$, we have in particular $\dim P_{n}=\#{\cal S}={{n+d}\choose d}=p_{n}$, and the matrix $\Theta_{{\cal S}}P=V:=V_{\leq n}({\cal S})$ is our Vandermonde matrix. Therefore, if we can find a sample set ${\cal S}$ such that the polynomial interpolation problem with respect to $\Theta_{{\cal S}}$ is poised for $P_{n}$, this would immediately imply that $|V|\neq 0$. Examples of such situations are provided in [chung1977, GascaMariano2000], and it turns out that there is a general sufficient condition which ensures poisedness.
Definition 8 ([chung1977]).
@@
Theorem 3 (Theorem 1 in [chung1977]).
Let $\mathcal{J}:=\{x_{1},...,x_{N}\}$ be a lattice of $N=p_{n}$ distinct nodes of $\mathbb{R}^{d}$. If $\mathcal{J}$ satisfies the GC condition as in Definition 8, then the polynomial interpolation problem with respect to $\mathcal{J}$ is poised for $P_{n}$.
@@
$P(x)=\sum_{i=1}^{N}y_{i}\prod_{j=1}^{n}{{G_{ij}(x)}\over{G_{ij}(x_{i})}}.$
Since $Y$ was arbitrary, uniqueness follows. ∎
Next, we recall the definition of $n$-th order principal lattice [nicolaides1972].
Definition 9.
@@
It was proven in [chung1977] Theorem 4, that $B(n,d)$ satisfies the GC condition which, by Theorem 3, guarantees poisedness [GascaMariano2000]. This is almost what we want, and indeed as we will show, the uniform grid contains a scaled principal lattice.
Lemma 4.
@@
$A(n,d):=n\cdot B(n,d)=\biggl{\{}x\in\mathbb{R}^{d}\Big{|}x=\sum_{i=0}^{d}\gamma_{i}X_{i},\sum_{i=0}^{d}\gamma_{i}=n,\gamma_{i}\in J_{n}\biggr{\}}\subset{\cal G}_{n}.$
@@
The proof is very similar to the one in [chung1977] for $B(n,d)$, and we provide it for completeness.
Let $\{\gamma_{k}:\mathbb{R}^{d}\to\mathbb{R}\}_{k=0,...,d}$ be the barycentric coordinate functions associated with the simplex $X_{0},...,X_{d}$. Define the following hyperplanes $H_{km}:=\{x|\gamma_{k}(x)=m\}$, $k=0,...,d$, $m=0,...,n$. These hyperplanes contain all the nodes in $A(n,d)$.
@@
Figure 1: Demonstration of the proof of Lemma 4.
Remark 5.
@@
Theorem 4.
Let ${\cal X}\subset\mathbb{R}^{d}$ be a finite set of $n$ distinct points in $\mathbb{R}^{d}$, and let $m=\mu({\cal X})$ be its discrete moment order. Let $\ell$ be the unique integer such that $p_{\ell-1}<n\leq p_{\ell}$. Then for every $N\geq\ell$ we have
@@
so the conclusion of Lemma 3 holds for ${\cal S}={\cal G}_{N}$.
@@
Using Lemma 4 and Theorems 2 and 3, we get
$\big{|}V_{\leq\ell}(A(\ell,d))\big{|}\neq 0$ $\implies\operatorname{rank}(V_{\leq\ell}(A(\ell,d)))={{\ell+d}\choose d}=p_{\ell}\geq n$
$A(\ell,d)\subset{\cal G}_{\ell}$ $\implies\operatorname{rank}(V_{\leq\ell}({\cal G}_{\ell}))\geq{{\ell+d}\choose d}\geq n.$
Now notice that since $\operatorname{rank}V_{\leq m}({\cal X})=n$ we must have $p_{m}\geq n$ and therefore $m\geq\ell$. This implies
$\operatorname{rank}(V_{\leq m}({\cal G}_{N}))\geq\operatorname{rank}(V_{\leq m}({\cal G}_{\ell})\geq\operatorname{rank}(V_{\leq\ell}({\cal G}_{\ell}))\geq n,$

### v2

@@ 3 Eigenvalues of smooth kernel matrices in the flat limit
The bounds of Theorem 3.1 are not necessarily tight. As we show in the next lemma, the exactness of the asymptotic orders depends on the analytic properties of the kernel $\mathcal{K}$ and the geometry of $\cal X$}, therefore additional assumptions ought to be made on both of those objects in order to obtain matching lower bounds.
Definition 3.3.
Let ${\cal X}\subset\mathbb{R}^{d}$. Consider the matrices $Q_{\perp,k-1}$ defined in the proof of Theorem 3.1. We define a sequence $\{\mathcal{M}_{k}\}_{k=0}^{m}$ of linear subspaces of $\mathbb{R}^{n}$ with $\dim\mathcal{M}_{k}=t_{k}$, $\mathcal{M}_{k}\subseteq\operatorname{range}(Q_{\perp,k-1})$ and $\mathcal{M}_{k}\perp\operatorname{range}(Q_{\perp,k})$. By a slight abuse of notation, we also denote by $\mathcal{M}_{k}\in\mathbb{R}^{n\times t_{k}}$ a matrix whose columns span $\mathcal{M}_{k}$. For consistency we set $\operatorname{range}(Q_{\perp,-1})\equiv\mathbb{R}^{n}$. Since $\sum_{i=0}^{m}t_{i}=n$, we have
$\mathcal{Q}_{\perp}:=\big(\mathcal{M}_{0}\;\mathcal{M}_{1}\;\dots\mathcal{M}_{m}\big)\in\mathbb{R}^{n\times n}.$
The matrix $\mathcal{Q}_{\perp}^{T}V_{\leq m}$ is upper triangluar. We denote its block-diagonal part by
@@
Lemma 3.4.
For $\mathcal{K}\in\mathcal{C^{(\infty,\infty)}}(\Omega)$ and ${\cal X}\subset\mathbb{R}^{d}$, let $\tilde{R}\in\mathbb{R}^{n\times p_{m}}$ be as in Definition 3.3. Define $C=\det\left(\tilde{R}W\tilde{R}^{T}\right)\in\mathbb{R}$ where $W=W^{\mathcal{K}}_{\leq m}\in\mathbb{R}^{p_{m}\times p_{m}}$ is the Wronskian matrix of $\mathcal{K}$ as in Definition 2.4. Then
Proof.
By [undef, Lemma 6.4] we have
@@
where the last equality follows from $\epsilon^{2{(m+1)}}E_{n}^{-1}=O(\epsilon)$ and $\tilde{R}W\tilde{R}^{T}\in\mathbb{R}^{n\times n}$ . This implies that $\epsilon^{-2\sum_{j=0}^{m}jt_{j}}\det(K_{\epsilon})=\det(\tilde{R}W\tilde{R}^{T}+O(\epsilon))=\det(\tilde{R}W\tilde{R}^{T})+O(\epsilon)$ (by Lemma (B.2) in the Appendix) with $C=\det(\tilde{R}W\tilde{R}^{T})$.
Corollary 3.5.
Suppose $\mathcal{K}$ is symmetric and analytic (in both variables) in the neighborhood of $(\mathbf{0},\mathbf{0})$. The scaling of the eigenvalues as given in Theorem 3.1 is exact, if and only if $\det(\tilde{R}W\tilde{R}^{T})\neq 0$. By ’exact scaling’ we mean that there exists a constant $\tilde{\lambda}_{k,j}\neq 0$ independent of $\epsilon$ such that $\lambda_{k,j}=\epsilon^{2k}(\tilde{\lambda}_{k,j}+O(\epsilon))$.
@@
By Theorem 3.1 and [undef, Theorem 2.9], the eigenvalues of $K_{\epsilon}({\cal X})$ satisfy
@@
where $t_{k}:=\operatorname{rank}(V_{\leq k})-\operatorname{rank}(V_{\leq k-1})$ and $m=\mu({\cal X})$, and $\tilde{\lambda}_{s,j}$ do not depend on $\epsilon$ (but may be zero). By Lemma 3.4 we have that
@@
[…]
@@ 4 Eigenvalues of the multidimensional Dirichlet kernel
In this section we consider the specific case of the multidimensional Dirichlet kernel, which is directly motivated by the problem of super-resolution as outlined in the Introduction. Let ${\cal S}\subset\mathbb{R}^{d}$ denote a finite set of frequencies, symmetric around the origin, where the spectral data (2) is available. We first derive a generic sufficient condition on $\cal S$} which ensures that the scaling order of Theorem 3.1 is exact. Later we show that this condition is satisfied by ${\cal S}={\cal G}_{N}$, the regular symmetric grid in $\mathbb{R}^{d}$ of side length $2N+1$, for large enough $N$.
Definition 4.1.
@@ 4.1 A nondegeneracy condition
Note that $\mathcal{D}_{{\cal S}}$ is a translation invariant, symmetric and analytic kernel. Recall the definition (8) of the Wronskian matrix for an arbitrary kernel. Set $W:=W^{\mathcal{D_{S}}}_{\leq m}$. By Corollary 3.5, in order to show tightness, it is sufficient to demonstrate that $C:=\det(\tilde{R}W\tilde{R}^{T})\neq 0$. In the following lemma, we show that this holds under a certain (more or less natural) non-degeneracy assumption on the set $\cal S$}.
Lemma 4.2.
For $W:=W_{\leq m}$ as defined above, $\det(\tilde{R}W\tilde{R}^{T})>0$ holds whenever $\operatorname{rank}(V_{\leq m}({\cal S}))\geq n$, where $V_{\leq m}({\cal S})=\big[{(\omega)}^{\alpha}\big]_{\omega\in{\cal S}}^{\alpha\in\mathbb{P}_{m}}$.
Remark 4.3.
@@
where $F:=\operatorname{diag}\left(\frac{i^{|\alpha|}}{\alpha!}\right)_{\alpha\in\mathbb{P}_{m}}$. Let $B:=FV^{T}_{\leq m}({\cal S})$, and consider $B=U\Sigma T^{*}$ the singular value decomposition of $B$. We therefore have $W=BB^{*}=U\Sigma^{2}U^{*}$, implying that $W$ is positive semidefinite (since $W$ is Hermitian and has non-negative eigenvalues). Since $\tilde{R}$ is full rank (recall Definition 3.3), it follows that $\tilde{R}W\tilde{R}^{T}$ is positive semidefinite as well, implying that $\det(\tilde{R}W\tilde{R}^{T})\geq 0$. To ensure $\det(\tilde{R}W\tilde{R}^{T})>0$, we require that $\tilde{R}W\tilde{R}^{T}$ be positive definite. The latter condition is satisfied when $\operatorname{rank}(\tilde{R}W\tilde{R}^{T})\geq n$, which is equivalent to $\operatorname{rank}(W)=\operatorname{rank}(V_{\leq m}({\cal S}))\geq n$. ∎
Remark 4.4.
For $d=1$, we have that m=n1m=n-1 and $p_{n-1}={n\choose 1}=n$, meaning that the condition $\operatorname{rank}(V_{\leq n-1}({\cal S}))\geq n$ in Lemma 4.2 is satisfied as long as the sampling set $\cal S$} contains at least $n$ distinct points.
By applying Corollary 3.5, we immediately obtain the following result.
Lemma 4.5.
@@ 4.2 Geometric characterization condition and the uniform grid
The rank condition in Lemma 4.2 is nontrivial to verify for a given set $\cal S$}. In this section we provide an explicit example of such a set, namely, the uniform grid ${\cal G}_{N}$ in $\mathbb{R}^{d}$. Our approach is a straightforward application of well-known results in multivariate polynomial interpolation, adapted to our setting.
We first recall some definitions and results from [undefak]. For a set $\Theta$, let $\Theta$ denote its cardinality. For a square matrix $M$, let $|M|$ denote its determinant, i.e. $|M|:=\det(M)$.
Definition 4.6 ([undefak], p.194).
Let $\mathcal{P}$ be a linear subspace of the polynomial ring $\Pi:=\mathbb{R}\left[x_{1},...,x_{d}\right]$. The polynomial interpolation problem with respect to a (finite) set of linearly independent functionals $\Theta\subset\Pi^{\prime}$ ($\Pi^{\prime}$ is the dual space of $\Pi$) is said to be poised for $\mathcal{P}$, if for any $Y\in\mathbb{R}^{\#\Theta}$ there exists a unique $f\in\mathcal{P}$ such that $\Theta f=Y$.
Theorem 4.7 ([undefak] p.194).
@@
For $\Theta_{{\cal S}}:=\{\delta_{\xi}\}_{\xi\in{\cal S}}$ the set of point evaluation functionals on $\cal S$}, and the standard monomial basis $P_{n}$ of polynomials of total degree $\leq n$, we have in particular $\dim P_{n}=\#{\cal S}={{n+d}\choose d}=p_{n}$, and the matrix $\Theta_{{\cal S}}P_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n}=V:=V_{\leq n}({\cal S})$ is our Vandermonde matrix. Therefore, if we can find a sample set $\cal S$} such that the polynomial interpolation problem with respect to $\Theta_{{\cal S}}$ is poised for $\mathcal{P}_{n}:=\text{span}\{P_{n}\}$, this would immediately imply that $|V|\neq 0$. Examples of such situations are provided in [undefi, undefq], and it turns out that there is a general sufficient condition which ensures poisedness.
Definition 4.8 ([undefi]).
@@
Theorem 4.9 (Theorem 1 in [undefi]).
Let $\mathcal{J}:=\{x_{1},...,x_{N}\}$ be a lattice of $N=p_{n}$ distinct nodes of $\mathbb{R}^{d}$. If $\mathcal{J}$ satisfies the GC condition as in Definition 4.8, then the polynomial interpolation problem with respect to $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Theta_{\mathcal{J}}$} is poised for $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}_{n}$}.
@@
$P(x)=\sum_{i=1}^{N}y_{i}\prod_{j=1}^{n}{{G_{ij}(x)}\over{G_{ij}(x_{i})}},$
where $G_{i1},\dots,G_{in}$ are distinct hyperplanes associated with $x_{i}\in\mathcal{J}$, and they satisfy the conditions stated in Definition 4.8. Since $Y$ was arbitrary, uniqueness follows. ∎
Next, we recall the definition of $n$-th order principal lattice [undefae].
Definition 4.10.
@@
It was proven in [undefi] Theorem 4, that $B(n,d)$ satisfies the GC condition which, by Theorem 4.9, guarantees poisedness [undefq]. This is almost what we want, and indeed as we will show, the uniform grid contains a scaled principal lattice.
Lemma 4.11.
@@
$A(n,d):=n\cdot B(n,d)=\biggl\{x\in\mathbb{R}^{d}\Big|x=\sum_{i=0}^{d}\gamma_{i}X_{i},\sum_{i=0}^{d}\gamma_{i}=n,\gamma_{i}\in J_{n}\biggr\}\subset{\cal G}_{n}.$ (12)
@@
The proof is very similar to the one in [undefi] for $B(n,d)$, and we provide it for completeness.
Let $\{\gamma_{k}:\mathbb{R}^{d}\to\mathbb{R}\}_{k=0,...,d}$ be the barycentric coordinate functions associated with the simplex $X_{0},...,X_{d}$. Define the following hyperplanes $H_{km}:=\{x|\gamma_{k}(x)=m\}$, k=0,,dk=0,...,d, m=0,,nm=0,...,n. These hyperplanes contain all the nodes in $A(n,d)$.
@@
Figure 1: Demonstration of the proof of Lemma 4.11.
Remark 4.12.
@@
For any $t\in\mathbb{R}^{d}$ and a set $A\subseteq\mathbb{R}^{d}$, we denote by $A_{t}$ the shifted set $A_{t}:=\{a+t:a\in A\}$. The following result is immediate.
Proposition 4.13.
If $A$ satisfies the GC condition of degree $n$, then $A_{t}$ satisfies the GC condition of degree $n$ for any $t\in\mathbb{R}^{d}$.
@@
Theorem 4.14.
Let ${\cal X}\subset\mathbb{R}^{d}$ be a finite set of $n$ distinct points in $\mathbb{R}^{d}$, and let $m=\mu({\cal X})$ be its discrete moment order. Let $\ell$ be the unique integer such that $p_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2\ell-2}<n\leq p_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2\ell-1}$. Then for every $N\geq\ell$ we have
@@
so the conclusion of Lemma 4.5 holds for ${\cal S}={\cal G}_{N}$.
@@
Let $s=-(\ell-1,\dots,\ell-1)^{T}\in\mathbb{R}^{d}$ and define $A_{s}(2\ell-1,d):=A(2\ell-1,d)+s$. By Lemma 4.11 and Proposition 4.13 we have that $A_{s}(2\ell-1,d)$ satisfies the GC condition of degree $2\ell-1$. Using Theorem 4.9 (or alternatively Corollary 4.18) we get that the polynomial interpolation problem with respect to $\Theta_{A_{s}(2\ell-1,d)}$ is poised for $\text{span}\{P_{2\ell-1}\}$, thus by Theorem 4.7 we get
[…]
$A_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2\ell-1},d)\subset{\cal G}_{\ell}$ $\implies\operatorname{rank}(V_{\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2\ell-1}}({\cal G}_{\ell}))\geq{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2\ell-1}+d}\choose d}\geq n.$
Now notice that since $\operatorname{rank}V_{\leq m}({\cal X})=n$ we must have $p_{m}\geq n$ and therefore $m\geq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2\ell-1}$. This implies
$\operatorname{rank}(V_{\leq m}({\cal G}_{N}))\geq\operatorname{rank}(V_{\leq m}({\cal G}_{\ell}))\geq\operatorname{rank}(V_{\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2\ell-1}}({\cal G}_{\ell}))\geq n,$
@@
Remark 4.15.
Theorem 4.14 can be reformulated without explicitly referring to the set $\mathcal{{\cal X}}$ and its discrete moment order $m=\mu(\mathcal{{\cal X}})$ by instead imposing the assumption that $p_{m}\geq n$ for given $m,n$.
We now introduce an alternative condition for poisedness, following the work [undefo].
Definition 4.16 ([undefo]).
A finite set $L\subset\mathbb{N}_{0}^{d}$ is a lower set if whenever $\mu\in L$ and $0\leq\alpha\leq\mu$ then $\alpha\in L$, where $\alpha\leq\beta$ means that $\alpha_{j}\leq\beta_{j}$ for $j=1,\dots,d$.
Theorem 4.17 (Theorem 1 in [undefo]).
Let $L$ be a lower set and $F_{L}:=\{x_{\alpha}:\alpha\in L\}$. Then the polynomial interpolation problem with respect to $\Theta_{F_{L}}$ is poised for $\mathcal{P}_{L}:=\text{span}\{x^{\alpha}:\alpha\in L\}$.
Corollary 4.18.
The polynomial interpolation problem with respect to $\Theta_{A_{s}(2\ell-1,d)}$ is poised for $\mathcal{P}_{2\ell-1}:=\text{span}\{P_{2\ell-1}\}$, where $P_{2\ell-1}$ is the standard monomial basis of polynomials of total degree $\leq 2\ell-1$ and $s:=-(\ell-1,\dots,\ell-1)^{T}$.
Proof.
It is easy to verify that the set $A(2\ell-1,d)$ is a lower set as in Definition 4.16. Now we show that $\mathcal{P}_{A(2\ell-1,d)}\equiv\mathcal{P}_{2\ell-1}$. By the definition of $A(2\ell-1,d)$ in (12), any $\alpha\in A(2\ell-1,d)$ has the form $(\alpha_{1},\dots,\alpha_{d})$ where $\alpha_{i}\in\{0,1,\dots,2\ell-1\}$ and $\sum_{i=0}^{d}\alpha_{i}=2\ell-1$ which is equivalent to $\sum_{i=1}^{d}\alpha_{i}\leq 2\ell-1$. Hence the set $\{x^{\alpha}:\alpha\in A(2\ell-1,d)\}$ is precisely the set of monomials of total degree $\leq d$. To complete the proof, let $F_{A(2\ell-1,d)}:=\{\alpha+s:\alpha\in A(2\ell-1,d)\}$. It is easy to see that $F_{A(2\ell-1,d)}\equiv A_{s}(2\ell-1,d)$, therefore, using Theorem 4.17, we get the desired result. ∎
Note that corollary 4.18, gives an alternative proof of the poisdness result in Theorem 4.14.